# Dial-In LLM: Human-Aligned LLM-in-the-loop Intent Clustering for Customer Service Dialogues

### v3

@@ Abstract
Discovering customer intentions in dialogue conversations is crucial for automated service agents. However, existing intent clustering methods often fail to align with human perceptions due to a heavy reliance on embedding distance metrics and a tendency to overlook underlying semantic structures. This paper proposes an LLM-in-the-loop (LLM-ITL) intent clustering framework, integrating the semantic understanding capabilities of LLMs into conventional clustering algorithms. Specifically, this paper (1) investigates the effectiveness of fine-tuned LLMs in semantic coherence evaluation and intent cluster naming, achieving over 95% accuracy aligned with human judgments; (2) designs an LLM-ITL framework that facilitates the iterative discovery of coherent intent clusters and the optimal number of clusters; and (3) proposes context-aware techniques tailored for customer service dialogue. As existing English benchmarks offer limited semantic diversity and intent groups, we introduce a comprehensive Chinese dialogue intent dataset, comprising over 100k real customer service calls and 1,507 human-annotated intent clusters. The proposed approaches significantly outperform LLM-guided baselines, achieving notable enhancements in clustering quality and lower computational cost. Combined with several best practices, our findings highlight the potential of LLM-in-the-loop techniques for scalable and human-aligned intent clustering.
@@
Mengze Hong1, Wailing Ng1, Chen Jason Zhang1 Yuanfeng Song2, Di Jiang2thanks: Corresponding author: dijiang@webank.com 1Hong Kong Polytechnic University 2AI Group, WeBank Co., Ltd
@@ 1 Introduction
Intent discovery is a crucial task in numerous NLP applications, including the design of AI-driven dialogue systems Hengst et al. (2024) and the learning of utterance patterns Ghosal et al. (2020). While intent clustering techniques are extensively studied to automatically identify thematic relationships within text corpora Gung et al. (2023), existing research primarily focuses on developing meaningful sentence representations and relies on embedding distance metrics for optimization Yin et al. (2021). This approach often overlooks the distinctive characteristics of textual information, such as linguistic patterns and semantic diversity (see Figure 1), and restricts human-aligned evaluation and validation of clustering quality Vinh et al. (2009). This issue is particularly pronounced in languages with rich semantics, such as Chinese, where two seemingly similar sentences can convey entirely different meanings, and vice versa Chen (1993).
With growing research interest in integrating Large Language Models (LLMs) into the problem-solving pipeline Hong et al. (2025), LLM-guided clustering techniques have emerged, demonstrating superior performance over traditional machine learning algorithms. While these methods effectively incorporate language understanding into the clustering process, they primarily focus on data preprocessing, querying LLMs for embedding refinement or data augmentation Zhang et al. (2023); Viswanathan et al. (2024). This approach represents a surface-level integration, potentially missing the benefits of LLMs in providing semantic-driven guidance to the clustering process and the quality of the final results.
In this paper, we introduce an LLM-in-the-loop (LLM-ITL) intent clustering method, designed to facilitate the iterative discovery of coherent intent clusters. Our approach adeptly utilizes intermediate clustering results by incorporating human-aligned LLM utilities, offering computationally efficient and human-interpretable intent clustering. The key contributions of this work are:
@@
    We introduce the largest Chinese dialogue intent clustering dataset, notable for its high semantic diversity and a significant inclusion of noisy, out-of-domain intents. This dataset is derived from over 100,000 real-world customer service call sessions, comprising 1,507 intent clusters annotated by human experts. It fills a crucial gap in evaluation benchmarks for complex intent clustering and classification.
@@
    We demonstrate the effectiveness of fine-tuned LLMs in assessing the semantic coherence of intent clusters and accurately providing intent labels across various sampling strategies. This highlights two effective and efficient utilities for designing LLM-in-the-loop solutions tailored for intent clustering.
@@
    We propose an LLM-in-the-loop clustering algorithm to derive an optimal set of clusters in scenarios where the number of clusters is unknown. This approach effectively leverages intermediate results generated during the clustering process, outperforming state-of-the-art baselines and providing practical benefits for downstream applications. Additionally, the discussion on data sampling and LLM-based crowdsourcing further improves performance.
@@ 2 Related Work
This study is grounded in two key areas of research: LLM-guided text clustering and intent clustering.
@@
The integration of LLMs into text clustering has become increasingly prominent since 2023. Zhang et al. (2023) introduced ClusterLLM, an innovative approach that leverages instruction-tuned LLMs like ChatGPT to refine sentence embedding spaces through pairwise preference questions, aligning clustering granularity with user preferences. Viswanathan et al. (2024) highlighted the enhancement of clustering quality by LLMs through feature improvement, the imposition of constraints during clustering, and post-correction processes. Additionally, Feng et al. (2024) proposed refining edge points with LLMs, which led to notable performance gains. Recently, Hong et al. (2025) conceptualized LLM-in-the-loop machine learning, classifying LLM integration into data-, model-, and task-centric strategies. This method mirrors a “human-in-the-loop” approach Chen et al. (2024), facilitating effective LLM integration in machine learning applications.
@@
Intent clustering is a fundamental step in intent induction Chandrakala et al. (2024) and the development of intent classifiers Gung et al. (2023). While semi-supervised Kumar et al. (2022) and deep learning methods Lin et al. (2020); Zhang et al. (2021) have significantly advanced intent clustering, recent progress in LLM-based systems has highlighted their potential practical advantages Liang et al. (2024). The IDAS method effectively employs in-context learning to generate descriptive utterance labels and enhance sentence embeddings De Raedt et al. (2023). However, challenges such as computational cost and robustness remain unresolved Song et al. (2023).
@@
Previous studies have primarily focused on enhancing input text representations and the direct use of LLMs for cluster modification, imposing high computational cost and uncertainty. In contrast, this paper implements human-aligned LLM utilities to create a LLM-in-the-loop intent clustering process with emphasis on the intermediate results. Additionally, this work tackles a more realistic scenario by releasing a complex dataset transcribed from real customer service calls, offering practical insights and contributing to more meaningful evaluations in future intent clustering.
Figure 2: Overview of proposed LLM-in-the-loop Dialogue Intent Clustering.
@@ 3.1 LLM Utilities
Inspired by human behaviors in perceiving texts and making comparisons based on semantic meaning rather than word-level similarity Peter W. Foltz and Landauer (1998), this paper proposes the semantic consistency within a cluster, known as semantic coherence, as a superior metric and optimization target. The coherence evaluation is formulated as a binary classification problem, leading toward intuitive evaluation and easing the training of the LLM evaluator (see Section 5.6 for different formulations). Good clusters consist of sentences focused on a specific topic, while Bad clusters contain inconsistent or ambiguous intentions (see Table 9 for examples). This task is particularly challenging for traditional machine learning classifiers due to a lack of semantic understanding Mimno et al. (2011), necessitating fine-tuned LLMs Gu et al. (2025) for robust evaluation (see Section 5.1 for fine-tuning details).
@@
Giving each cluster a concise and meaningful name is essential for many downstream applications Pattnaik et al. (2024). In this paper, we introduce a novel naming convention, "Action-Objective," which is particularly effective for capturing dialogue intents that are typically topic-oriented (e.g., insurance, loan) and involve distinct actions (e.g., inquire, confirm). Examples of human annotations are shown in Table 10, and a comparison of different naming conventions is presented in Table 11, demonstrating the superiority of the proposed convention in capturing customer intention with clear goal-oriented expression.
@@ 3.2 LLM-in-the-loop Iterative Intent Clustering with Coherence Evaluation
3:$\mathcal{E}\leftarrow\{f_{\text{emb}}(s)\mid s\in\mathcal{S}\}$
4:$\mathcal{S}^{(0)}\leftarrow\mathcal{S}$, $\mathcal{C}\leftarrow\emptyset$, $t\leftarrow 0$
@@
$\mathcal{C}_{n_{i}}^{(t)}=F(\mathcal{E}^{(t)},n_{i}),$
where $F$ is a clustering function (e.g., K-means clustering), and $\mathcal{E}^{(t)}=\{\mathbf{e}_{s}\mid s\in\mathcal{S}^{(t)}\}$ representing the sentence embeddings. This results in $N$ distinct cluster assignments.
While the number of clusters is often known with benchmark evaluation, it is challenging but necessary to implement mechanisms for determining the optimal number of clusters in a noisy text corpus. Here, we propose a local search heuristic that maximizes the “good/bad” ratio at each iteration. The optimal $n_{t}^{*}$ is selected as:
where $n_{t}$ denotes the best cluster number at iteration $t$. This enables the automatic discovery of suitable cluster numbers in a step-by-step manner. The search space $\bm{N}$ should be chosen carefully to balance accuracy and efficiency; thus, we propose a search space pruning strategy, as illustrated in Section 5.6 for more details.
@@ 3.3 Post-Correction with LLM-Generated Intent Labels
Preliminary results in Table 12 suggest that multiple clusters may capture similar intents due to diminishing size of unassigned sentences $\mathcal{S}^{(t)}$ in later iterations, leading to smaller and less representative clusters. The embedding distances between sentences within clusters limit the natural consolidation of clusters with similar intents but different expressions Khan et al. (2020), necessitating a post-correction step to merge semantically aligned clusters. In contrast to previous methods that rely on direct LLM queries for validating individual cluster assignments Viswanathan et al. (2024); Feng et al. (2024), which are computationally expensive and inconsistent, we propose a context-aware approach, leveraging LLMs’ naming utility to robustly merge clusters based on intent labels.
@@
with $\pi_{m}=1/K$ as uniform mixture weights. An edge is retained only if this probability exceeds a predefined threshold $\tau$ ($\tau=0.7$), minimizing inappropriate merges by accounting for uncertainty.
@@ 3.4 Context-Aware Role Separation with LLM-Generated Intent Labels
From a practical perspective, dialogue datasets are often accompanied by domain-specific features or use-case scenarios (e.g., customer service calls, group discussions). This motivates the incorporation of contextual information into the clustering process. The customer service dialogues typically involve only the customer and the service agent Lin et al. (2022). The classification task that assigns a sentence to its associated role is relatively simple with labeled training data. To provide an unsupervised solution, the sentence roles are naturally determined based on the LLM- intent label. By heuristics, sentences within clusters labeled as “inquire-” or similar actions are mostly sent from the customer, and vice versa. A two-step approach is proposed: initially, intermediate clustering results are obtained using previous approach, denoted as $C_{\text{inter}}=\{C_{1},C_{2},\ldots,C_{K}\}$, and these results are divided based on intent labels $l_{k}$, forming two groups with distinct roles, $R_{\text{customer}}$ and $R_{\text{agent}}$, where:
$\begin{array}[]{c}R_{\text{customer}}=\{s\in S\mid l_{k}(s)\in\{\text{"inquire-"},\dots\}\},\\ R_{\text{agent}}=\{s\in S\mid l_{k}(s)\notin\{\text{"answer-"},\dots\}\}.\end{array}$
In the second step, the sentences corresponding to each role are clustered again, denoted as $C^{\prime}_{\text{customer}}=\text{Cluster}(R_{\text{customer}})$ and $C^{\prime}_{\text{agent}}=\text{Cluster}(R_{\text{agent}})$, to achieve more refined outcomes, allowing for a clear separation of customer intent from service agent intent. Finally, the two groups of clusters are merged together, serving as the final or intermediate result to be further processed.
@@ 4.1 Dataset
The proposed dataset contains 1,507 high-quality intent clusters manually annotated from over 100,000 realistic customer service calls, comprising 55,085 distinct sentences with an average length of 17 Chinese characters per sentence (see Appendix A for annotation details) [^1] The full dataset and fine-tuned models will be released upon acceptance. Sample dataset can be found at GitHub.. Among these intents, 885 are identified as domain-specific queries (e.g., inquire-insurance), while 622 are categorized as out-of-domain (e.g., provide-location), representing general queries that commonly arise in a customer service interaction. This dataset is the first Chinese benchmark for intent clustering and the largest of its kind compared to the existing English benchmark datasets shown in Table 1. It presents several new challenges, including the lexical sparsity in short-text sentences, the dependence on contextual knowledge, and the difficulties in balancing accuracy with computational efficiency due to the excessively large intent size. The semantic diversity, calculated based on the average cosine distance between individual sentences and the cluster centroid Casanueva et al. (2022), depicts the complexity of this dataset and motivates the development of semantic-guided techniques.
@@
Table 1: Comparison of proposed dataset with existing intent clustering datasets.
Category Method NMI #Good
Baselines K-Means 0.7899 94.8%
GMMs 0.7903 91.1%
Hierarchical* 0.8001 94.9%
LLM-Guided Clustering Baselines ClusterLLM Zhang et al. (2023) 0.7284 91.2%
IDAS De Raedt et al. (2023) 0.8109 93.6%
Keyphrase Viswanathan et al. (2024) 0.8371 94.5%
LLMEdgeRefine Feng et al. (2024) 0.7411 87.2%
Proposed Method (context-free) LLM-ITL (random) 0.8202 97.7%
LLM-ITL (convex) 0.8208 97.8%
LLM-ITL + keyphrase 0.8378 96.4%
Proposed Method (context-aware) LLM-ITL + merge 0.8420 97.8%
LLM-ITL + role 0.8679 97.2%
LLM-ITL + role + merge 0.8826 97.6%
Table 2: Performance of baselines and proposed methods in dialogue intent clustering. The best results are bolded.
{CJK}
UTF8gkai
LLM qwen7b qwen14b baichuan2-7b chatglm3-6b
Accuracy 96.25% 97.50% 89.17% 95.83%
{CJK}
UTF8gkai
LLM qwen7b qwen14b baichuan2-7b chatglm3-6b
Accuracy 92.8% 94.3% 94.3% 94.4%
Table 4: Performance of fine-tuned LLMs in naming intent clusters.

### v4

@@ Abstract
Discovering customer intentions is crucial for automated service agents, yet existing intent clustering methods often fall short due to their reliance on embedding distance metrics and neglect of underlying semantic structures. To address these limitations, we propose an LLM-in-the-loop (LLM-ITL) intent clustering framework, integrating the language understanding capabilities of LLMs into conventional clustering algorithms. Specifically, this paper (1) examines the effectiveness of fine-tuned LLMs in semantic coherence evaluation and intent cluster naming, achieving over 95% accuracy aligned with human judgments; (2) designs an LLM-ITL framework that facilitates the iterative discovery of coherent intent clusters and the optimal number of clusters; and (3) introduces context-aware techniques tailored for customer service dialogue. Since existing English benchmarks lack sufficient semantic diversity and intent coverage, we further present a comprehensive Chinese dialogue intent dataset comprising over 100k real customer service calls with 1,507 human-annotated clusters. The proposed approaches significantly outperform LLM-guided baselines, achieving notable improvements in clustering quality, cost efficiency, and downstream applications. Combined with several best practices, our findings highlight the prominence of LLM-in-the-loop techniques for scalable dialogue data mining.
@@
Mengze Hong1thanks: Work was partially done during internship at WeBank., Wailing Ng1, Chen Jason Zhang1, Yuanfeng Song2, Di Jiang1thanks: Corresponding Author 1Hong Kong Polytechnic University, 2AI Group, WeBank Co., Ltd
@@ 1 Introduction
Intent discovery is a crucial task in NLP applications, such as dialogue system design Hengst et al. (2024), information retrieval Jiang et al. (2016a), and utterance pattern analysis Ghosal et al. (2020). While intent clustering techniques are extensively studied to automatically identify thematic relationships within text corpora Gung et al. (2023), existing research primarily focuses on developing meaningful sentence representations and relies on embedding distance metrics for optimization Yin et al. (2021); Gao et al. (2025). This approach often overlooks the distinctive characteristics of textual information, such as linguistic patterns and semantic diversity (see Figure 1), and thus restricts human-aligned evaluation and validation of clustering performance Vinh et al. (2009). This issue is particularly pronounced in languages with rich semantics, such as Chinese, where two seemingly similar sentences can convey entirely different meanings Chen (1993); Jiang et al. (2016b).
With growing research interest in integrating large language models (LLMs) into the problem-solving pipeline Hong et al. (2025b), LLM-guided clustering techniques have emerged, demonstrating superior performance over traditional machine learning algorithms Zhang et al. (2023). While these methods effectively incorporate language understanding into the clustering process, they primarily focus on data preprocessing, querying LLMs for embedding refinement or data augmentation Viswanathan et al. (2024). This approach represents a surface-level integration, potentially missing the benefits of LLMs to contribute semantic-driven guidance within the clustering process.
In this paper, we introduce an LLM-in-the-loop (LLM-ITL) intent clustering framework, designed to facilitate the iterative discovery of coherent intent clusters from semantically diverse, large-scale dialogue datasets. Our approach effectively leverages intermediate clustering results by incorporating human-aligned LLM utilities, enabling computationally efficient and human-interpretable intent clustering. The key contributions of this work are:
@@
    We present the largest Chinese dialogue intent clustering dataset, derived from over 100,000 real-world customer service calls across the banking, telecommunication, and insurance domains. The data is annotated into 1,507 intent clusters with high semantic diversity and a substantial inclusion of noisy, out-of-domain queries, reflecting realistic and complex customer interactions.
@@
    We demonstrate the effectiveness of fine-tuned small LLMs in assessing the semantic coherence of intent clusters and providing accurate intent labels across various sampling strategies, offering cost-efficient utilities for designing LLM-in-the-loop solutions.
@@
    We propose an LLM-in-the-loop intent clustering framework that effectively combines the strengths of LLMs and conventional clustering algorithms. This approach outperforms state-of-the-art baselines and excels in downstream applications with 18.46% performance gain. Furthermore, discussions on data sampling and LLM-based crowdsourcing validate best practices for real-world deployment.
Figure 2: Overview of the proposed LLM-in-the-loop framework for dialogue intent clustering.
@@ 2 Related Work
The integration of LLMs into text clustering has become increasingly prominent since 2023. Zhang et al. (2023) introduced ClusterLLM, an innovative approach that leverages instruction-tuned LLMs like ChatGPT to refine sentence embedding spaces through pairwise preference questions, aligning clustering granularity with user preferences. Viswanathan et al. (2024) highlighted the enhancement of clustering quality by LLMs through feature improvement, the imposition of constraints during clustering, and post-correction processes. Additionally, Feng et al. (2024) proposed refining edge points with LLMs, which led to notable performance gains.
Recently, Hong et al. (2025b) introduced the concept of LLM-in-the-loop machine learning, categorizing integration strategies into data-, model-, and task-level approaches. This paradigm parallels the human-in-the-loop framework Chen et al. (2024), enabling LLMs to replicate human expertise and thereby enhance conventional problem-solving workflows in a cost-efficient and flexible manner. Within this taxonomy, existing work on LLM-guided clustering has primarily focused on data-centric aspects, with relatively limited exploration of modeling and task-solving aspects. This gap constrains a broader understanding of the potential of LLMs in tackling long-standing challenges in text clustering, such as improving cluster interpretability and moderating the clustering process Tan et al. (1999); Jiang et al. (2021).
@@
Intent clustering extends conventional text clustering by incorporating contextual cues and domain knowledge to uncover meaningful user intentions Allahyari et al. (2017); Hong et al. (2024). It serves as a fundamental step in intent induction Chandrakala et al. (2024) and in preparing data for training intent classifiers Gung et al. (2023). While conceptually related to topic modeling Jiang et al. (2015, 2023), which seeks to uncover latent semantic themes, intent clustering tackles the more challenging task of differentiating texts that may appear lexically similar but convey contextually distinct meanings, thereby uncovering the underlying communicative goals behind sentences Carberry and Flowers (1988); Qu et al. (2018). This finer granularity makes intent clustering more demanding as a data mining task and a critical component in applications such as search engines Jiang et al. (2013, 2016a) and dialogue systems Qin et al. (2023); Hong et al. (2025c).
While semi-supervised approaches Kumar et al. (2022) and deep learning methods Lin et al. (2020); Zhang et al. (2021) have substantially advanced intent clustering, recent attempts in LLM-based systems highlight their practical advantages Liang et al. (2024). For instance, the IDAS method leverages in-context learning to generate descriptive utterance labels and improve sentence embeddings De Raedt et al. (2023). However, challenges such as high computational cost and limited model robustness remain unresolved Song et al. (2023), and existing methods are evaluated based on small-scale datasets with limited complexity Casanueva et al. (2020), leading to an inadequate understanding of their effectiveness in real-world applications.
@@
Previous studies have primarily focused on enhancing input text representations and the direct use of LLMs for cluster modification, imposing high computational cost and uncertainty. In contrast, this paper implements human-aligned LLM utilities to create an LLM-in-the-loop framework with emphasis on the intermediate clustering results. Additionally, this work releases a complex dataset from real customer service call transcriptions, enabling more practical insights and meaningful evaluations for future intent clustering.
@@ 3.1 LLM Utilities
Inspired by human behaviors in perceiving texts and making comparisons based on semantic meaning rather than surface-level work similarity Peter W. Foltz and Landauer (1998), this paper proposes semantic coherence as a more effective metric and optimization objective, measuring the semantic consistency within a cluster. The coherence evaluation is formulated as a binary classification problem, enabling intuitive interpretation and simplifying the training of the LLM evaluator (see Section 5.6 for alternative formulations). Good clusters consist of sentences focused on a specific topic, whereas Bad clusters contain inconsistent or ambiguous intentions (see Table 11 for examples). This task is particularly challenging for traditional machine learning classifiers due to their lack of semantic understanding Mimno et al. (2011), thus necessitating fine-tuned LLMs for robust evaluation Gu et al. (2025).
@@
Giving each cluster a concise and meaningful name is essential for many downstream applications Pattnaik et al. (2024); Luo et al. (2024). In this paper, we introduce a novel naming convention, ”Action-Objective,” which is particularly effective for capturing dialogue intents that are typically topic-oriented (e.g., insurance, loan) and involve distinct actions (e.g., inquire, confirm). Examples of human annotations are shown in Table 12, and a comparison of different naming conventions is presented in Table 13 to demonstrate the effectiveness of the proposed approach.
@@ 3.2 LLM-in-the-loop Iterative Intent Clustering with Coherence Evaluation
$\mathcal{C}_{n_{i}}^{(t)}=F(\mathcal{E}^{(t)},n_{i}),$
where $F$ is a clustering function (e.g., K-means clustering), and $\mathcal{E}^{(t)}=\{\mathbf{e}_{s}\mid s\in\mathcal{S}^{(t)}\}$ represents the sentence embeddings. This results in $|N|$ distinct cluster assignments as the initial outcome.
While the number of clusters is often known in benchmark evaluations, determining the optimal number in a noisy text corpus is both challenging and essential. Here, we propose a local search heuristic that maximizes the “good/bad” ratio at each iteration. The optimal $n_{t}^{*}$ is given by:
representing the best cluster number at iteration $t$. This approach enables the automatic discovery of suitable cluster numbers in a step-by-step manner. The search space $\bm{N}$ should be selected carefully to balance accuracy and efficiency, and a search space pruning strategy is proposed in Section 5.6 to enhance the searching process.
@@
[…]
@@ 3.3 Post-Correction with LLM-Generated Intent Labels
Preliminary results in Table 14 suggest that, in later iterations, the diminishing size of the unassigned sentence set $\mathcal{S}^{(t)}$ may lead to the formation of multiple clusters capturing similar intents, resulting in smaller and less representative clusters. The embedding distances between sentences within clusters limit the natural consolidation of clusters with similar intents but different expressions Khan et al. (2020), necessitating a post-correction step to merge semantically aligned clusters. Previous methods typically address this by issuing direct LLM queries to validate individual cluster assignments Viswanathan et al. (2024); Feng et al. (2024), an approach that is both computationally expensive and prone to inconsistency. In contrast, we propose a context-aware approach, leveraging LLMs’ naming utility to robustly merge clusters based on their generated intent labels.
@@
with $\pi_{m}=1/K$ as uniform mixture weights. An edge is retained only if this probability exceeds a predefined threshold $\tau$ (e.g., $\tau=0.7$), minimizing inappropriate merges by accounting for uncertainty.
@@ 3.4 Context-Aware Role Separation with LLM-Generated Intent Labels
From a practical perspective, dialogues are often accompanied by domain-specific features or use-case scenarios (e.g., customer service calls, group discussions). This motivates the incorporation of contextual information into the clustering process Ding et al. (2025). In particular, the customer service dialogues typically involve only the customer and the service agent Lin et al. (2022). The classification task that assigns a sentence to its associated role is relatively simple with labeled training data. To provide an unsupervised solution, we propose that the sentence roles can be naturally determined based on the LLM-generated intent label.
By heuristics, sentences within clusters labeled as “inquire-” or similar actions are mostly sent from the customer, and vice versa. A two-step approach is proposed: initially, intermediate clustering results are obtained using previous approach, denoted as $C_{\text{inter}}=\{C_{1},C_{2},\ldots,C_{K}\}$, and these results are divided based on intent labels $l_{k}$, forming two groups with distinct roles, $R_{\text{customer}}$ and $R_{\text{agent}}$, where:
$\begin{array}[]{c}R_{\text{customer}}=\{s\in S\mid l_{k}(s)\in\{\text{"inquire-"},\dots\}\},\\ R_{\text{agent}}=\{s\in S\mid l_{k}(s)\in\{\text{"answer-"},\dots\}\}.\end{array}$
In the second step, the sentences corresponding to each role are clustered again, denoted as $C^{\prime}_{\text{customer}}=\text{Cluster}(R_{\text{customer}})$ and $C^{\prime}_{\text{agent}}=\text{Cluster}(R_{\text{agent}})$. Finally, the resulting customer and agent clusters are merged to produce a refined clustering that maintains a clear separation of intents by role, serving either as the final output or as an improved intermediate stage for further processing.
@@ 4.1 Dataset
The proposed dataset contains 1,507 high-quality intent clusters manually annotated from over 100,000 realistic customer service calls, comprising 55,085 distinct sentences with an average length of 17 Chinese characters per sentence (see Appendix A for annotation details) [^1] Data is available at GitHub repository.. Among these intents, 885 are identified as domain-specific (e.g., inquire-insurance), primarily concentrated within the banking, telecommunications, and insurance industries, with a focus on the Chinese context Hong et al. (2025a). The remaining 622 clusters are categorized as out-of-domain (e.g., provide-location), representing general queries that commonly arise in customer service interactions.
Compared with existing intent clustering benchmarks such as BANK77 Casanueva et al. (2020) and CLINC150 Larson et al. (2019), the proposed dataset is the first Chinese benchmark for customer service intent clustering and the largest of its kind in both the number of sentences and the number of clusters (see Table 1). It presents several new challenges, including the lexical sparsity in short-text sentences, the dependence on contextual knowledge, and the difficulties in balancing accuracy with computational efficiency due to the excessively large intent size. The semantic diversity, calculated based on the average cosine distance between individual sentences and the cluster centroid Casanueva et al. (2022), depicts the complexity of this dataset and motivates the development of semantic-guided approaches.
@@
Table 1: Comparison of the proposed customer service intent clustering dataset with existing benchmarks.